OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

I have a basic understanding for these types of problems, I am just having some difficulty with this one.

OpenStudy (zarkon):

\[7 u^2 - 14 u + 2 = -5\]

OpenStudy (anonymous):

Yes just like that

OpenStudy (anonymous):

For other problems I have factored, Im just not sure what to do with this one. Should I subtract -2 from both sides? Should I factor out the 7 from both sin terms?

OpenStudy (zarkon):

\[7 u^2 - 14 u + 7=0\]

OpenStudy (anonymous):

Okay thats what I did the first time. I then factored out 7 to be left with \[7(\sin^2x-2sinx+1)=0\]

OpenStudy (zarkon):

\[7(\sin(x)-1)^2=0\]

OpenStudy (anonymous):

That was the point where I ended. I didnt know where to take it from there.

OpenStudy (zarkon):

\[\Large\Rightarrow \sin(x)-1=0\]

OpenStudy (anonymous):

Which we would need to isolate the sin to get \[sinx=1\]

OpenStudy (anonymous):

What happens to the 7 though?

OpenStudy (zarkon):

divide both sides by 7

OpenStudy (anonymous):

so sinx=1/7? The only thing that confuses me about that, is that there is no sine or y value on the unit circle that is 1/7

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[7(\sin(x)-1)^2=0\] \[\Rightarrow (\sin(x)-1)^2=0/7=0\]

OpenStudy (zarkon):

so you need just to find \[\sin(x)=1\]

OpenStudy (anonymous):

how does 0/7=0?

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

Nevermind I get it now

OpenStudy (zarkon):

zero divided by any nonzero number is zero

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

That would be like multiplication. 0 times any number is zero.